For a 60 kg mass, if the friction force is 40 N, the magnitude of the applied force required to accelerate the object at 4.0m/s^2 is ____x10^2 N (Give your answer with the correct number of Significant digits and do not include units)

Note: Let East be the positive direction


1
Expert's answer
2021-07-08T09:57:57-0400

According to Newton's second law:


"ma=F-f,\\
F=ma+f=2.8\cdot10^2\text{ N}"